Two opposite sides of a square are increased by twenty-five percent while the other two are decreased by forty percent. Find the percent of decrease in the area of the square.
(A) 15% (B) 25% (C) 30% (D) 32.5% (E) none of these

Answer by stanbon(75887)

Two opposite sides of a square are increased by twenty-five percent while the other two are decreased by forty percent. Find the percent of decrease in the area of the square.
Area of original square: s^2 where s is the length of a side
Area of new rectangle:
((5/4)s)((6/10)s) = (3/4)s^2
Quantity of decrease = s^2 - (3/4)s^2 = (1/4) s^2
% of decrease = [(1/4)s^2]/s^2 = 1/4 = 25%
=================
Cheers,
Stan H.
(A) 15% (B) 25% (C) 30% (D) 32.5% (E) none of these